OpenStudy (anonymous):

Solve the equation for y x=3y+6

OpenStudy (acxbox22):

x=3y+6 x-6=3y ------- 3 x/3 -2=y

OpenStudy (anonymous):

So would it be y=x-2 ?

OpenStudy (acxbox22):

no, it would be \[y=\frac{ x }{ 3 }-2\]

OpenStudy (anonymous):

Ok thank you

OpenStudy (acxbox22):

np